# PPFPL: Cross-silo Privacy-preserving Federated Prototype Learning Against Data Poisoning Attacks on Non-IID Data

### v4

@@ VII-B6 Stability of PPFPL
It is observed from Fig. 4 and 5 that the performance of PPFPL exhibits some fluctuations. Therefore, it is important to evaluate the stability of PPFPL over multiple randomized experiments. To verify the stability of PPFPL, we conduct experiments using the MNIST, FMNIST, and CIFAR10 datasets, with 20% of clients being malicious. The data distribution is Non-IID (i.e., $\textit{Avg}=3,\textit{Std}=2$). Each experiment is repeated 10 times. The utilization of boxplots in Fig. 8 provides a clear visualization of the stability of PPFPL. The results show that PPFPL’s performance fluctuates less and remains stable against data poisoning attacks. The stability of PPFPL proves its reliability in real-world scenarios.
VII-B7 Scalability of PPFPL
Federated learning is often deployed with an uncertain number of clients, and scalability directly determines whether PPFPL can maintain stable performance in environments with different numbers of clients. To evaluate the scalability of PPFPL, we conducted experiments on the MNIST, FMNIST, and CIFAR10 datasets with 20% malicious clients, and the number of clients set to 20, 25, 30, 35, and 40. The data distribution is Non-IID (i.e., $\textit{Avg}=3,\textit{Std}=1$). Fig. 9 shows the scalability of PPFPL against data poisoning attacks. It is observed that the performance of PPFPL drops by 0.24% on MNIST and 0.42% on CIFAR10 when the number of participants is increased by 100%, which is normal in federated learning. Thus, PPFPL shows excellent scalability against data poisoning attacks with data heterogeneity.
@@ VII-B7 Scalability of PPFPL
VII-B8 Impact of Local Iterations
We investigate the effect of the number of local iterations on the test average accuracy for PPFPL. We conduct experiments on MNIST, FMNIST, and CIFAR10 datasets, where 20% of malicious clients launched feature attacks. The number of local iterations are 1, 5, 10, 15, 20 and 25. The results are shown in Fig. 10. We observe that when the number of local iterations is 1, the changes in the local model are very small, resulting in slow training and relatively low accuracy for the same number of communication rounds. When the number of local iterations is 5, PPFPL has the highest test average accuracy on the MNIST dataset. However, on CIFAR10, the optimal number of local iterations should be set to 15. In addition, when the number of local iterations is too high, the test average accuracy decreases, which is due to the overfitting of the local model.
((a)) MNIST.
((b)) FMNIST.
((c)) CIFAR10.
Figure 10: The test average accuracy on three datasets with different number of local iterations under feature attacks.
VII-B9 Efficiency Evaluation
@@
To evaluate the efficiency of PPFPL, we measure the number of parameters submitted by clients, as shown in Fig. 11. We can observe that since the number of parameters submitted by PBFL and ShieldFL depends on the model architecture, their clients submit the same amount of parameters. Furthermore, the number of parameters submitted by clients in PPFPL is much lower than in PBFL under the same model architecture. This is because the number of prototype parameters in PPFPL depends on the compression of the feature extractor’s output, rather than the model architecture. Thus, PPFPL reduces the number of parameters submitted to two servers, thereby reducing the privacy computation overheads and communication overheads.
TABLE III: The cost time of ciphertext operation using CNN model under PBFL, ShieldFL, and PPFPL.
@@
In addition, ciphertext operations are the primary factor affecting the efficiency of our framework. We evaluate the time cost of PPFPL, PBFL, and ShieldFL on ciphertext operations. Note that PBFL is a federated learning scheme based on CKKS, while ShieldFL is built upon the two-door HE. For each client, we use the time spent on encryption and decryption for each iteration as a metric. As shown in Table III, PPFPL has advantages in encryption and decryption compared to PBFL and ShieldFL. This is because the number of prototype parameters submitted by clients is lower in PPFPL, which greatly improves the computational efficiency and reduces the computational cost.
VII-C Complexity Analysis
TABLE IV: Computation Overload and Communication Overload of client.
@@ VII-C Complexity Analysis
We analyze the computational overload and communication overload for each client in PPFPL and compared it with similar schemes [16][18][15]. The results are shown in Table IV. The computational overload for the clients in PPFPL comprises three components: local model training, prototype generation, and encryption. Formally, the computational overhead is expressed as $\mathcal{O}(T_{tr})+\mathcal{O}(T_{pro})+\mathcal{O}(pT_{ch})$, where $p$ denotes the prototype parameters, $T_{ch}$ denotes the time overload of encryption, $T_{tr}$ denotes the time overhead of local model training, and $T_{pro}$ denotes the time overhead of prototype generation. Notably, the overhead of prototype generation is significantly lower than that of local model training, i.e., $T_{tr}\gg T_{pro}$. Other similar schemes [16][18][15] adopt homomorphic encryption on gradients, their encryption overhead is $\mathcal{O}(gT_{ch})$, where $g$ denotes the gradient parameters. Since $g>p$, the overload satisfies $\mathcal{O}(gT_{ch})>\mathcal{O}(pT_{ch})$. Hence, the relationship $\mathcal{O}(T_{tr})+\mathcal{O}(gT_{mul})+\mathcal{O}(gT_{ch})\textgreater\mathcal{O}(T_{tr})+\mathcal{O}(gT_{ch})\textgreater\mathcal{O}(T_{tr})+\mathcal{O}(T_{pro})+\mathcal{O}(pT_{ch})$ holds. Consequently, PPFPL imposes a reasonable computational overhead on clients.
In addition, the client communication overhead is $\mathcal{O}(nP_{h})$ in PPFPL, where $P_{h}$ denotes the communication complexity of a number. Compared with [16][18][15], the relationship $\mathcal{O}(gP_{h})\textgreater\mathcal{O}(pP_{h})$ holds.
VIII Concludes
In this paper, we propose PPFPL, a privacy-preserving federated prototype learning framework, which improves performance while resisting data poisoning attacks in poisoned Non-IID data. Our work is the first to introduce prototype learning into PPFL to address privacy threats and data poisoning attacks in across-silo FL. As for how to implement lightweight privacy-preserving FL to resist data poisoning attacks on Non-IID data, we leave them to future exploration.

### v5

@@ 7.2.7 Stability of PPFPL
It is observed from Fig. 4 and 5 that the performance of PPFPL exhibits some fluctuations. Therefore, it is important to evaluate the stability of PPFPL over multiple randomized experiments. To verify the stability of PPFPL, we conduct experiments on MNIST, FMNIST, and CIFAR10 datasets, with 20% of clients being malicious. The data distribution follows the Non-IID setting (Avg=3, Std=2). Each experiment is repeated 10 times. The boxplots in Fig. 7 provides a clear visualization of the stability of PPFPL. The results show that PPFPL’s performance fluctuates less and remains stable against data poisoning attacks. The stability of PPFPL proves its reliability in real-world scenarios.
7.2.8 Scalability of PPFPL
Federated learning is often deployed on an unknown number of clients. Scalability directly determines whether PPFPL can maintain stable performance under different client scales. To evaluate the scalability of PPFPL, we conduct experiments on MNIST, FMNIST, and CIFAR10 datasets with 20% malicious clients, varying the number of clients as 20, 30, 40, 80, and 100. The data distribution is Non-IID (i.e., $\textit{Avg}=3,\textit{Std}=1$) setting. Fig. 8 shows the scalability of PPFPL against data poisoning attacks. It is observed that PPFPL maintains stable performance when the number of clients is 40. However, the performance of PPFPL degrades significantly when the number increases to 100. For example, the accuracy drops by 1.72% on MNIST and 14.11% on CIFAR10. The main reason is that PPFPL follows a personalized training paradigm in which each client relies on its own data for training. As the number of clients increases, the amount of data allocated to each client decreases. The small amount of data in each client makes it difficult to obtain a high-quality model through local training. This shows that the performance of PPFPL is closely related to the number of clients and ultimately depends on on the data available to each client. Therefore, in large-scale networks, each client needs to have sufficient data to fully exploit the advantages of PPFPL. Under our experimental conditions, the data volumes of MNIST, FMNIST, and CIFAR10 datasets can reasonably support up to about 40 clients, within which PPFPL demonstrates good scalability.
@@ 7.2.8 Scalability of PPFPL
7.2.9 Ablation Study
To comprehensively evaluate the effectiveness of the secure aggregation protocol, we conduct ablation experiments. Within the PPFPL framework, we remove the normalization verification module (denoted as “w/o NV”, where “w/o” stands for “without”) and the secure aggregation protocol (denoted as “w/o SAP”). We conduct experiments on CIFAR10 dataset, where 30% of malicious clients launched attacks. The experimental results are shown in TABLE 5. Under both attack settings, PPFPL’s performance degrades regardless of which component is removed, demonstrating the effectiveness of normalization verification and the secure aggregation protocol. Notably, in the absence of malicious clients, removing the secure aggregation protocol actually led to optimal model performance. This phenomenon is attributed to the increase in the number of benign training samples, which enhances the overall FL performance.
Table 5: Performance impact of different components in PPFPL.
  Schemes
  Data
  Distribution
  Test Average Accuracy (%)
  Feature   Label   No
  PPFPL   Avg=3, Avg=1   82.31   82.50   83.07
  Avg=3, Avg=2   83.16   83.42   83.29
  w/o NV   Avg=3, Avg=1   80.26   80.29   83.20
  Avg=3, Avg=2   80.74   80.61   83.34
  w/o SAP   Avg=3, Avg=1   77.69   77.24   83.53
  Avg=3, Avg=2   78.03   78.53   83.64
Note: “Feature” indicates feature attacks, “Label” indicates label attacks, and “No” indicates the absence of malicious clients. In addition, “NV” indicates the normalization verification module, “SAP” indicates the secure aggregation protocol. Best rusults are in blod.
7.2.10 Efficiency Evaluation
@@
To evaluate the efficiency of PPFPL, we measure the number of parameters submitted by clients, as reported in Fig. 9. We observe that both PBFL and ShieldFL require clients to submit a fixed number of parameters determined by the model architecture, resulting in identical parameter volumes across all clients. In contrast, the number of parameters submitted by clients in PPFPL is significantly lower under the same model architecture. This is because the number of prototype parameters in PPFPL is determined by the output of the feature extractor rather than the model architecture itself. Consequently, PPFPL effectively reduces the parameter volume transmitted to the servers, thereby decreasing both computational and communication overhead.
Table 6: The cost time of ciphertext operation using CNN model under PBFL, ShieldFL, and PPFPL.
@@
Furthermore, ciphertext operations are the primary factor affecting the efficiency of our framework. We evaluate the time cost of ciphertext operations for PPFPL, PBFL, and ShieldFL. Notably, PBFL employs the CKKS, while ShieldFL is built upon two-door HE. We measure the time spent on encryption and decryption per round for each client as our evaluation metric. As shown in Table 6, PPFPL’s performance in both encryption and decryption is lower than that of PBFL and ShieldFL. This advantage stems from the fact that PPFPL requires fewer prototype parameters submitted by each client, significantly reducing the computational cost of the ciphertext operations.
7.3 Complexity Analysis
Table 7: Computation overload and communication overload of client.
@@ 7.3 Complexity Analysis
We analyze the computational and communication overload for each client in PPFPL and compared it with similar schemes [16][19][15], as shown in TABLE 7. The computational overload for clients in PPFPL consists of three components: local model training, prototype generation, and encryption. Formally, the computational overhead can be expressed as $\mathcal{O}(T_{tr})+\mathcal{O}(T_{pro})+\mathcal{O}(pT_{ch})$, where $p$ is the number of prototype parameters, $T_{ch}$ denotes the encryption time overhead, $T_{tr}$ denotes the time overhead for local model training, and $T_{pro}$ denotes the time overhead of prototype generation. Notably, the overhead of prototype generation is significantly lower than that of local model training, i.e., $T_{tr}\gg T_{pro}$. Other comparable schemes [16][19][15] adopt homomorphic encryption on gradients, resulting in an encryption overhead $\mathcal{O}(gT_{ch})$, where $g$ denotes the number of gradient parameters. Since the relation $g\textgreater p$ holds, the overload relation satisfies $\mathcal{O}(gT_{ch})\textgreater\mathcal{O}(pT_{ch})$. Hence, the relation of ciphertext operations is $\mathcal{O}(gT_{mul})+\mathcal{O}(gT_{ch})\textgreater\mathcal{O}(gT_{ch})\textgreater\mathcal{O}(pT_{ch})$. In addition, the client communication overhead in PPFPL is $\mathcal{O}(nP_{h})$, where $P_{h}$ denotes the communication complexity per parameter. Compared with the schemes in [16][19][15], the inequality $\mathcal{O}(gP_{h})\textgreater\mathcal{O}(pP_{h})$ holds, demonstrating the efficiency advantage of PPFPL.
8 Conclusion
This paper proposes PPFPL, a privacy-preserving federated prototype learning framework that effectively enhances federated learning performance in poisoned Non-IID data environments. Through comprehensively theoretical analysis and experimental validation, we demonstrate the significant advantages of PPFPL compared existing FL defense schemes. The framework provides robust security guarantees for distributed computing scenarios, with important practical applications in privacy-sensitive domains such as finance and healthcare.
However, our current framework has certain limitations. First, frequent interactions in the two-server architecture increase system failures caused by communication latency. Second, existing experimental datasets do not adequately capture the complexity of real-world cross-silo scenarios, making it difficult to effectively evaluate prototypes generated by benign clients under varying data quality conditions. This limitation may result in unfair evaluation outcomes in practical settings. Therefore, our future work will focus on two directions: (i) designing more efficient two-server interaction protocols while ensuring privacy and security. (ii) constructing benchmark datasets that capture the characteristics of cross-silo scenarios to evaluate FL system performance.